OpenStudy (anonymous):

What is the slope of the line through the points (–2, –1) and (8, –3)?

OpenStudy (just_gymnast_me):

do you know the slope formula?

OpenStudy (just_gymnast_me):

Its \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

it's -1/5

OpenStudy (just_gymnast_me):

then you should take the "y value" from the second parentheses and subtract by the first y value

OpenStudy (just_gymnast_me):

-3 - -1 = -2

OpenStudy (just_gymnast_me):

then do the same with the x values

OpenStudy (just_gymnast_me):

8 - -2 = 10

OpenStudy (just_gymnast_me):

-2/10 simplified is -1/5

OpenStudy (anonymous):

thank you u help me with a few more

OpenStudy (anonymous):

please?(:

OpenStudy (just_gymnast_me):

im not very good with these but i can try, i have to go very soon so only one more.

OpenStudy (anonymous):

Which equation is in standard form and represents a line with slope 5 through the point (0, 2)?

OpenStudy (just_gymnast_me):

whats are the choices?

OpenStudy (anonymous):

y=5x+2 5x+y=2 y-2=5x -5x+y=2

OpenStudy (just_gymnast_me):

first one

OpenStudy (anonymous):

thank u